# VividDreamer: Invariant Score Distillation For Hyper-Realistic Text-to-3D Generation

### v1

@@ 1 Introduction
This paper considers text-to-3D content generation [16, 49, 37, 34]. It is a fundamental task in computer vision with the potential to benefit various downstream applications, such as AR/VR, and gaming. Benefiting from large-scale and high-quality text-image pairs, there recently emerges a new 3D generation method, called 2D lifting methods. The main idea of these methods is to utilize text-to-image generative models [46, 44, 10, 39, 21] trained on large-scale text-image pairs to guide the 3D creation. Early methods like DreamFields [15] and CLIP-forge [43] use the CLIP [39, 21] as guidance. Recently, inspired by the tremendous success of diffusion models [46, 44, 10] in the realm of image generation and conditional image generation [33, 41, 40], DreamFusion [37] introduces a loss based on probabilistic density distillation [35], called Score Distillation Sampling (SDS). They utilize pre-trained diffusion priors to optimize 3D representations [31, 32, 19] by steering their rendered images toward regions with higher probability density, conditioned on the text. This technology has sparked an abundance of subsequent research efforts. For example, Liu et al. [25] propose Zero123, which introduces SDS into the single-image reconstruction task [38, 24, 27]. However, despite its widespread impact, the 3D objects generated by SDS frequently suffer from over-saturation and over-smoothing problems[49, 1, 54].
@@
In this paper, we take an in-depth analysis of these issues. We decompose the SDS into two components, one is the reconstruction term: $\epsilon_{\phi}(z_{t};y,t)-\epsilon$, and the other is the classifier-free guidance term: $\epsilon_{\phi}(z_{t};y,t)-\epsilon_{\phi}(z_{t};\varnothing,t)$. We use each term individually as loss for text-to-image generation, and find two problems: 1) using the reconstruction term alone not only lacks semantic alignment functionality but also deteriorates the originally high-quality images. 2) using the classifier-free guidance term is capable of aligning images with textual prompts, but suffers from over-saturation (as shown in Fig. 2). SDS avoids problem 1 by assigning a large guidance scale (100.0) to the classifier-free guidance term, but this exacerbates problem 2.
@@
To overcome these problems, we analyze the causes of problem 1 and focus on optimizing the reconstruction term. We find the function of this item is actually to reconstruct the noise image in single-step, and this leads to problem 1 from two aspects: 1) At the beginning of the generation, reconstructing semantic-free images hinders semantic alignment. 2) When the image is aligned with the condition, the inconsistency of the single-step reconstruction destroys the image and causes over-smoothing. To address these, we propose the Invariant Score Distillation (ISD) method, which replaces the reconstruction item with the invariant score term. This term no longer reconstructs the noisy image, but aligns the noisy image to the previous time step state. Specifically, we derive a general sampling prior from the diffusion sampling formula, and follow the setting of DDIM [44] to remove the random noise to obtain our final invariant score prior. This brings two benefits: Firstly, the invariant score prior can be continuously refined during the generation process to provide more details for the results while avoiding errors caused by single-step reconstruction, avoiding over-smoothing and details lost problems. Secondly, the introduction of invariant score prior allows us to use a conventional guidance scale (7.5) for text-to-3D generation, avoiding the over-saturation problem. We also conduct extensive experiments to verify the effectiveness of our method. The results show that our method can generate realistic and detailed 3D objects in single-stage optimization. Compared with existing methods, our method shows strong competitiveness both qualitatively and quantitatively.
@@ 2.1 Text to image Generation
Recently, significant progress has been made in text-to-image generation [40, 39, 55, 21]. Among these, diffusion models [44, 46, 10, 17, 29] are particularly notable for their outstanding performance. These models, once trained, are capable of transforming Gaussian noise into high-definition, realistic samples through multiple iterations. To enable conditional image generation, Dhariwal et al. [7] introduces classifier guidance. This approach requires training an additional noisy image classifier, thereby needing extra gradient guidance during sampling, and limiting generated image variety. To address these issues, Ho et al. propose classifier-free guidance [11], eliminating the need for an explicit classifier. This method trains the diffusion model in both conditioned and unconditioned modes, it modifies the model by Eq. 1
@@ 2.2 Text to 3D Generation
Text-to-3D generation refers to creating 3D assets based on text descriptions. Current works can be categorized into two main types: (i) Native 3D methods that directly train a new diffusion model with 3D data [34, 12, 16, 20, 53] and (ii) 2D lifting methods that use the pre-trained 2D text-to-image diffusion models to guide the 3D assets generation [37, 48, 49, 54, 14, 52, 18, 22, 50, 23, 3]. Upon completion of pre-training, the former can rapidly generate 3D-consistent assets. However, due to the scarcity and expense of 3D datasets [2, 6, 5], the performance and diversity of current 3D diffusion models are significantly constrained.
On the other hand, 2D lifting methods leverage models pre-trained on rich 2D text-image datasets to guide the generation of 3D assets [32, 31, 19]. Dreamfusion [37] proposes Score Distillation Sampling (SDS) to utilize diffusion models for guiding the generation of 3D assets. Despite the popularity of this approach, empirical observations reveal that SDS often suffers from over-saturation and over-smoothing issues [49, 54, 14, 52, 23, 3]. To address these challenges, Magic3D [23] adopts a two-stage method by combining NeRF [31] with mesh fine-tuning, Fantasia3D [3] takes a unique approach by decoupling geometry and appearance, facilitating the generation of explicit meshes and textures. Additionally, some methods refine the loss function to promote the quality of created 3D models. HiFA [54] introduces an image-level loss term based on the original SDS. ProlificDreamer [49] proposes VSD, it replaces the random Gaussian noise with a learned score function of noisy rendered images conditioned on the camera pose $c$. NFSD [18] decomposes components of SDS and provides experimental improvements to the loss function, while CSD [52] analyzes the components of the SDS and directly uses the classifier score term as guidance. Recently, with the rise of the new rendering technology 3DGS [19], a new batch of excellent works has emerged [22, 47, 51, 4]. This paper also focuses on the improvement of the loss function, unlike others, we introduce a invariant score term inspired by DDIM. This term ultimately enables us to achieve high-quality 3D asset generation results in a single-stage optimization process.
@@ 3.1 Score Distillation Sampling
Score Distillation Sampling [37] is a novel technique in text-to-3D generation. It utilizes a pre-trained text-to-image diffusion model $\phi$ to guide the 3D representation parameterized by $\theta$. Specifically, for a given camera pose $\pi$, we denote $x=g(\theta;\pi)$ as the image rendered from a differentiable rendering function $g$. SDS introduces a loss function to ensure that the images obtained from any viewpoint via the differentiable rendering function $g$ are aligned with the textual prompt $y$. The form of this loss function is as follows:
$\nabla_{\theta}\mathcal{L}_{SDS}=\mathbf{E}_{t,\epsilon,\pi}[w(t)(\epsilon_{\phi}(z_{t};y,t))\frac{\partial x}{\partial\theta}]$ (2)
where $w(t)$ is a weighting function, and $z_{t}$ represents a noisy version of $x$, obtained through $z_{t}=\sqrt{\alpha_{t}}x+\sqrt{1-\alpha_{t}}\epsilon$. In practice, DreamFusion [37] discovered that subtracting a Gaussian noise $\epsilon$ from the loss enables better performance. This change does not alter the objective since $\mathbf{E}_{t,\epsilon,\pi}(-\epsilon)=0$. On the other hand, SDS employs Classifier-free guidance [11] in practice and assigns a large guidance weight (100.0). Therefore, in reality, the $\epsilon_{\phi}(z_{t};y,t)$ in Eq. 2 is $\epsilon_{\phi}(z_{t};y,t)+w(\epsilon_{\phi}(z_{t};y,t)-\epsilon_{\phi}(z_{t};\varnothing,t))$, where $w$ represents the guidance scale. Combine them, the final form of SDS is:
@@
In this form, SDS consists of two parts. The first part is the original $\epsilon_{\phi}(z_{t};y,t)-\epsilon$, which is proportional to $\nabla_{z_{t}}\log p(z_{t}|y)$. Its role is to restore $z_{t}$ back to its pre-noised state $x$ according to the condition $y$. The second part is introduce by Classifier-free guidance [11], $\epsilon_{\phi}(z_{t};y,t)-\epsilon_{\phi}(z_{t};\varnothing,t)$, which is proportional to $\nabla_{z_{t}}\log p(z_{t}|y)-\nabla_{z_{t}}\log p(z_{t})$. According to Bayes’ rule, this term can further be proportional to $\nabla_{z_{t}}\log p(y|z_{t})$. It act as an implicit noisy image classifier, whose function is to determine whether $z_{t}$ meets the condition $y$.
@@ 4 Methodology
In this section, we introduce our proposed Invariant Score Distillation method for high-fidelity text-to-3D generation. Specifically, in Sec. 4.1, we decompose the SDS into two components: the classifier-free guidance term and the reconstruction term, and re-examine the role of each term in the generation process. We find the reconstruction term hinders the quality of the generation. To solve these problems, we introduce Invariant Score Distillation In Sec. 4.2, which replaces the reconstruction term with an invariant score term. Moreover, we also summarize the differences and connections between our method and other methods.
@@ 4.1 Anatomy of SDS
Based on the analysis in Sec. 3.1, we seek to explore why SDS suffers from over-smoothing and over-saturation. To this end, we conduct an experiment on 2D text-to-image to investigate the role of each term in the generation process of SDS. Specifically, following our previous analysis, we divide the SDS loss into two components: one is $\epsilon_{\phi}(z_{t};y,t)-\epsilon_{\phi}(z_{t};\varnothing,t)$, which we call the classifier-free guidance term, and the other part is $\epsilon_{\phi}(z_{t};y,t)-\epsilon$, which we refer to as the reconstruction term. We use a randomly generated noise and a high-quality image sampled from Stable Diffusion [40] as our initialization images. Then we separately use each term as the loss function for image generation. As demonstrated in Fig. 2, we find that using $\epsilon_{\phi}(z_{t};y,t)-\epsilon_{\phi}(z_{t};\varnothing,t)$ alone can generate images that align with the textual condition $y$, but it suffers from issues of over-saturation and distortion, this cannot be avoided even when using a high-quality image as initialization. In practice, SDS assigns a large guidance scale (100.0) to this term, thus exacerbating the problem of over-saturation.
@@
In this form, we find that $\epsilon_{\phi}(z_{t};y,t)-\epsilon$ is actually related to the $x_{0}^{t}$ predicted at time step $t$, meaning its function is to reconstruct $z_{t}$ back to its pre-noised state $x$. However, this presents two problems. One is the single-step reconstruction capability of the diffusion models is not strong [42, 45, 26], they can not reconstruct the image well when $t$ is large. This is why using $\epsilon_{\phi}(z_{t};y,t)-\epsilon$ alone causes an originally high-quality image to become noise (as shown in the third column of Fig. 2). Another issue is that when the original image $x$ has noise, reconstructing it back to its original state still retains the noise (as shown in the third column in Fig. 4). This directly leads to the following problems in SDS: at the beginning (when $x$ lacks semantic information), $\epsilon_{\phi}(z_{t};y,t)-\epsilon$ inhibits $\epsilon_{\phi}(z_{t};y,t)-\epsilon_{\phi}(z_{t};\varnothing,t)$ from aligning $x$ with the text condition $y$, and in the later stage of SDS (when $x$ already meets the text condition $y$), $\epsilon_{\phi}(z_{t};y,t)-\epsilon$ degrades $x$ into noise. Therefore, these two issues ultimately lead to over-smoothing and loss of detail in SDS optimization.
@@ 4.2 Invariant Score Distillation
To overcome the aforementioned problems, i.e., the over-smoothing and details lost problems brought by the reconstruction term and the over-saturation problem brought by the classifier-free guidance term. We introduce an invariant score term derived from the DDIM sampling process to replace the original reconstruction term. This term no longer reconstructs the noisy image to its original state in a single step but aligns the noisy image to the previous time step state, avoiding the error of single-step reconstruction.
@@
where $\lambda(t)$ and $w(t)$ is weighting function, we keep the value of $w(t)$ consistent with SDS. In addition to the advantages discussed above, we found that ISD also allows us to assign a normal weight (7,5) to the classifier terms, avoiding the problem of over-saturation. In Fig. 3, we present an overview of our framework. Our framework can optimize a Neural Radiance Field (NeRF) [31, 32] or 3D Gaussian Splatting [19] using 2D text-to-image diffusion models. In Fig. 1 and Fig. 5, we show our proposed ISD can generate high-fidelity and detailed 3D models in a single-stage optimization.
@@ 4.3 Connection with other methods
Compared with NFSD. NFSD [18] decomposes the loss function into three terms and obtains the final form by omitting the denoising term $\delta_{N}=\epsilon_{\phi}(z_{t};y,t)-\epsilon$:
@@ 5.2 Text-to-3D Generation
We present some representative results using ISD in Fig. 1 and 5, including both text-to-3DGS and text-to-NeRF (bottom two lines in Fig. 5). In general, ISD can generate 3D assets aligned well with textual prompts and possess a realistic and detailed appearance. For text-to-3DGS, we show that ISD can generate highly intricate details. For text-to-NeRF, we also show that ISD can effectively generate long text descriptions of content.

### v2

@@ 1 Introduction
This paper considers text-to-3D content generation [17, 51, 38, 35, 53, 3]. It is a fundamental task in computer vision with the potential to benefit various downstream applications. Benefiting from large-scale text-image pairs, there recently emerges a new 3D generation method, called 2D lifting methods. The main idea of these methods is to utilize text-to-image generative models [47, 45, 11, 40, 22] to guide the 3D creation. Early methods like DreamFields [16] and CLIP-forge [44] use the CLIP [40, 22] as guidance. Recently, inspired by the tremendous success of diffusion models [47, 45, 11] in the realm of image generation and conditional image generation [34, 42, 41, 54], DreamFusion [38] introduces a loss based on probabilistic density distillation [36], called Score Distillation Sampling (SDS). They utilize pre-trained diffusion priors to optimize 3D representations [32, 33, 20] by steering their rendered images toward regions with higher probability density, conditioned on the text. This technology has sparked an abundance of subsequent research efforts. For example, Liu et al. [26] propose Zero123, which introduces SDS into the single-image reconstruction task [39, 25, 28]. However, despite its widespread impact, the 3D objects generated by SDS frequently suffer from over-saturation and over-smoothing problems[51, 1, 61].
@@
In this paper, we take an in-depth analysis of these issues. We decompose the SDS into two components, one is the reconstruction term: $\epsilon_{\phi}(z_{t};y,t)-\epsilon$, and the other is the classifier-free guidance term: $\epsilon_{\phi}(z_{t};y,t)-\epsilon_{\phi}(z_{t};\varnothing,t)$. We use each term individually as loss for text-to-image generation, and find two problems: 1) using the reconstruction term alone not only lacks semantic alignment functionality but also deteriorates the originally high-quality images. 2) using the classifier-free guidance term alone suffers from serious over-saturation (as shown in Fig. 2). SDS avoids problem 1 by assigning a large guidance scale (100.0) to the classifier-free guidance term, but this exacerbates problem 2.
@@
To overcome these problems, we further analyze the causes of problem 1 and focus on optimizing the reconstruction term. We find the function of this item is actually to reconstruct the noise image in single-step, and this leads to problem 1 from two aspects: 1) At the beginning of the generation, reconstructing semantic-free images hinders semantic alignment. 2) When the image is aligned with the condition, the inconsistency of the single-step reconstruction destroys the image and causes over-smoothing. To address these, we propose the Invariant Score Distillation (ISD) method, which replaces the reconstruction item with the invariant score term. This term no longer reconstructs the noisy image, but aligns the noisy image to the previous time step state. This brings two benefits: Firstly, the invariant score prior can be continuously refined during the generation process to provide more details for the results while avoiding errors caused by single-step reconstruction, avoiding over-smoothing and details lost problems. Secondly, the introduction of invariant score prior allows us to use a conventional guidance scale (7.5) for text-to-3D generation, avoiding the over-saturation problem. We also conduct extensive experiments to verify the effectiveness of our method. The results show that our method can generate realistic and detailed 3D objects. Compared with existing methods, our method shows strong competitiveness both qualitatively and quantitatively.
@@ 2.1 Text to image Generation
Recently, significant progress has been made in text-to-image generation [41, 40, 62, 22, 59]. Among these, diffusion models [45, 47, 11, 18, 30] are particularly notable for their outstanding performance. These models, once trained, are capable of transforming Gaussian noise into high-definition, realistic samples through multiple iterations. To enable conditional image generation, Dhariwal et al. [8] introduces classifier guidance. This approach requires training an additional noisy image classifier, thereby needing extra gradient guidance during sampling, and limiting generated image variety. To address these issues, Ho et al. proposes classifier-free guidance [12], eliminating the need for an explicit classifier. This method trains the diffusion model in both conditioned and unconditioned modes, it modifies the model by Eq. 1:
@@ 2.2 Text to 3D Generation
Text-to-3D generation refers to creating 3D assets based on text descriptions. Current works can be categorized into two main types: (i) Native 3D methods that directly train a new diffusion model with 3D data [35, 13, 17, 21, 58] and (ii) 2D lifting methods that use the pre-trained 2D text-to-image diffusion models to guide the 3D assets generation [38, 49, 51, 61, 15, 57, 19, 23, 52, 24, 4, 60, 55, 50]. Upon completion of pre-training, the former can rapidly generate 3D-consistent assets. However, due to the scarcity and expense of 3D datasets [2, 7, 6], the performance and diversity of current 3D diffusion models are significantly constrained.
2D lifting methods leverage models pre-trained on rich 2D text-image datasets to guide the generation of 3D assets [33, 32, 20]. Dreamfusion [38] proposes Score Distillation Sampling (SDS) to utilize diffusion models for guiding the generation of 3D assets. Despite the popularity of this approach, empirical observations reveal that SDS often suffers from over-saturation and over-smoothing issues [51, 61, 15, 57, 24, 4]. To address these challenges, Magic3D [24] adopts a two-stage method by combining NeRF [32] with mesh fine-tuning, Fantasia3D [4] takes a unique approach by decoupling geometry and appearance, facilitating the generation of explicit meshes and textures. Additionally, some methods refine the loss function to promote the quality of created 3D models. HiFA [61] introduces an image-level loss term based on the original SDS. ProlificDreamer [51] proposes VSD, it replaces the random Gaussian noise with a learned score function of noisy rendered images conditioned on the camera pose $c$. NFSD [19] decomposes components of SDS and provides experimental improvements to the loss function, while CSD [57] analyzes the components of the SDS and directly uses the classifier score term as guidance. Recently, with the rise of the new rendering technology 3DGS [20], a new batch of excellent works has emerged [23, 48, 56, 5].
@@ 3.1 Score Distillation Sampling
Score Distillation Sampling (SDS) [38] utilizes a pre-trained text-to-image diffusion model $\phi$ to guide the 3D representation parameterized by $\theta$. Specifically, for a given camera pose $\pi$, we denote $x=g(\theta;\pi)$ as the image rendered from a differentiable rendering function $g$. SDS ensures that the images obtained from any viewpoint via the differentiable rendering function $g$ are aligned with the textual prompt $y$. The form of SDS loss is as follows:
$\nabla_{\theta}\mathcal{L}_{SDS}=\mathbf{E}_{t,\epsilon,\pi}[w(t)(\epsilon_{\phi}(z_{t};y,t)-\epsilon)\frac{\partial x}{\partial\theta}]$ (2)
where $w(t)$ is a weighting function, and $z_{t}$ represents a noisy version of $x$. In practice, SDS employs Classifier-free guidance [12], therefore, the $\epsilon_{\phi}(z_{t};y,t)$ in Eq. 2 is $\epsilon_{\phi}(z_{t};y,t)+w(\epsilon_{\phi}(z_{t};y,t)-\epsilon_{\phi}(z_{t};\varnothing,t))$, where $w$ represents the guidance scale. Combine them, the final form of SDS is:
@@
In this form, SDS consists of two parts. The first part is the original $\epsilon_{\phi}(z_{t};y,t)-\epsilon$, which is proportional to $\nabla_{z_{t}}\log p(z_{t}|y)$. Its role is to restore $z_{t}$ back to its pre-noised state $x$ according to the condition $y$. The second part is introduced by Classifier-free guidance [12], $\epsilon_{\phi}(z_{t};y,t)-\epsilon_{\phi}(z_{t};\varnothing,t)$, which is proportional to $\nabla_{z_{t}}\log p(z_{t}|y)-\nabla_{z_{t}}\log p(z_{t})$. According to Bayes’ rule, this term can further be proportional to $\nabla_{z_{t}}\log p(y|z_{t})$. It acts as an implicit noisy image classifier.
@@ 4.1 Anatomy of SDS
Based on the analysis in Sec. 3.1, we seek to explore why SDS suffers from over-smoothing and over-saturation. To this end, we conduct an experiment to investigate the role of each term in the generation process of SDS. Specifically, following our previous analysis, we divide the SDS loss into two components: one is the classifier-free guidance term $\epsilon_{\phi}(z_{t};y,t)-\epsilon_{\phi}(z_{t};\varnothing,t)$, and the other is the reconstruction term $\epsilon_{\phi}(z_{t};y,t)-\epsilon$. We use a randomly generated noise and a high-quality image sampled from Stable Diffusion [41] as our initialization images, respectively. Then we use each term as the loss function for image generation.
Figure 3: In order to show the relationship between over-saturation and guidance scale $w$ in SDS, we use different guidance scales for text-to-image generation. We find that as $w$ increases, the over-saturation of the generated images gradually becomes serious, while as $w$ decreases, the images gradually become over-smoothing.
As demonstrated in Fig. 2, we find that although using $\epsilon_{\phi}(z_{t};y,t)-\epsilon_{\phi}(z_{t};\varnothing,t)$ alone can generate images aligned with the textual condition $y$, it suffers from serious over-saturation and distortion, this cannot be avoided even when using a high-quality image as initialization. To further investigate the relationship between over-saturation and the classifier-free guidance term, we conduct another experiment. As shown in Fig. 3, we find that as $w$ increases, the images generated by SDS gradually become full of details (such as the texture of the horse, the details of the astronaut), but the over-saturation problem is also gradually amplified. Therefore, we speculate that the over-saturation phenomenon in SDS comes from the larger guidance scale. This phenomenon is also reasonable because Stable Diffusion [41] has not used classifier-free guidance term alone for generating during pre-training. Therefore, this item itself cannot generate real samples, and amplifying its weight have a negative impact on generated results.
@@
In this form, we find that $\epsilon_{\phi}(z_{t};y,t)-\epsilon$ is actually related to the $x_{0}^{t}$ predicted at time step $t$, meaning its function is to reconstruct $z_{t}$ back to its pre-noised state $x$. However, this presents two problems. One is the single-step reconstruction error caused by diffusion models [43, 46, 27], they can not reconstruct the image well when $t$ is large. This is why using $\epsilon_{\phi}(z_{t};y,t)-\epsilon$ alone causes an originally high-quality image to become noise (as shown in the third column of Fig. 2). Another is that when the original image $x$ has noise, reconstructing it back to its original state still retains the noise (as shown in the third column in Fig. 5). This directly leads to the following problems in SDS: at the beginning (when $x$ lacks semantic information), $\epsilon_{\phi}(z_{t};y,t)-\epsilon$ inhibits $\epsilon_{\phi}(z_{t};y,t)-\epsilon_{\phi}(z_{t};\varnothing,t)$ from aligning $x$ with the text condition, and in the later stage of SDS (when $x$ already meets the text condition), $\epsilon_{\phi}(z_{t};y,t)-\epsilon$ degrades $x$ into noise. Therefore, these two issues ultimately lead to over-smoothing and loss of detail in SDS optimization.
@@ 4.2 Invariant Score Distillation
To overcome the aforementioned problems, i.e., the over-smoothing and details lost problems brought by the reconstruction term and the over-saturation problem brought by the large classifier-free guidance scale. We introduce an invariant score term derived from the DDIM sampling process to replace the original reconstruction term. This term no longer reconstructs the noisy image to its original state in a single step but aligns the noisy image to the previous time step state, avoiding the error of single-step reconstruction.
@@
where $\lambda(t)$ and $w(t)$ is weighting function, we keep the value of $w(t)$ consistent with SDS. In addition to the advantages discussed above, we found that ISD also allows us to assign a normal weight (7,5) to the classifier terms, thus avoiding the problem of over-saturation. In Fig. 4, we present an overview of our framework. Our framework can optimize a Neural Radiance Field (NeRF) [32, 33] or 3D Gaussian Splatting [20] using 2D text-to-image diffusion models. In Fig. 1 and Fig. 6, we show our proposed ISD can generate high-fidelity and detailed 3D models in a single-stage optimization.
@@ 4.3 Connection with other methods
Compared with NFSD. NFSD [19] decomposes the loss function into three terms and omits the denoising term $\delta_{N}=\epsilon_{\phi}(z_{t};y,t)-\epsilon$. The final form is:
@@ 5.2 Text-to-3D Generation
We present some representative results using ISD in Fig. 1 and 6, including both text-to-3DGS and text-to-NeRF (bottom two lines in Fig. 6). In general, ISD can generate 3D assets aligned well with textual prompts and possess a realistic and detailed appearance. For text-to-3DGS, we show that ISD can generate highly intricate details. For text-to-NeRF, we also show that ISD can effectively generate long text descriptions of content.